kaitlynmills13:

Divide: 8xy^2-14y^2/2y

jhonyy9:

this wan being in this form : \[(8xy^2 -14y^2)/2y \] ?

jhonyy9:

than yes - in numerator you need factorizing out 2y and simplify it by denominator hope helped

kendra4640:

this is what I found out K = h-(c/r) b = (4/3)*(j-4) 3;7;1;1;1;0.056 2;12;3;12;3;1 0.15;3.20